# Text-to-Decision Agent: Learning Generalist Policies from Natural Language Supervision

### v2

@@ 3.4 Generalist Policy Learning
A widely adopted manner to tackle generalization in RL domains is to condition the policy on some representation that can capture task-relevant information, such as context-based meta-learning [23, 24] and prompt-based methods [8, 10]. Under this unified framework, we formulate the generalist agent as a task-conditioned policy model. There exists a true variable that represents the identity of the underlying task. Since we do not have access to this ground-truth information, we use a latent representation $h$ to approximate that variable. Then, a base policy $\pi(\cdot,h)$ can be shared across tasks by conditioning on the task representation as
$\pi_{k}(a\mid s)=\pi(a\mid s,h_{k}),~{}~{}~{}\forall k.$ (7)
Previous studies usually approximate the task representation by acquiring expert decision data [8] or exploring the unseen task first [9]. This restricted form of supervision limits their universality and practicability since expensive high-quality samples or explorations are needed to specify any new task.
Motivated by the recent success of language models, we study a promising alternative that leverages a much broader source of supervision. Natural language is a flexible representation for transferring a variety of human ideas and intentions. Learning from its supervision could not only enhance the representation power but also enable efficient zero-shot transfer. We map the user-provided textual description of the task $l$ to a latent embedding $\psi(l)$ via the fine-tuned text encoder in Sec. 3.3, and use this contextual embedding to approximate the task representation as $h\approx\psi(l)$. Then, the task-agnostic generalist policy is approximated by
$\pi(a\mid s,h_{k})\approx\pi\left(a\mid s,\psi(l_{k})\right),~{}~{}~{}\forall k.$ (8)
@@ 3.5 Scalale Implementations
where $c$ denotes the timestep in the forward diffusion process, and $c=0$ corresponds to the unperturbed data.  [^1] Decision diffuser diffuses over only states and trains another inverse dynamics model to generate actions. In contrast, we only keep a unified diffusion model to i) ensure simplicity with a single end-to-end model in a multi-task setting, and ii) unlock the potential of diffusion models at scale, rather than depending on additional MLPs to generate part of the information. To enable a generalist policy with natural language supervision, we include the fine-tuned text embedding $\psi(l)$ as an additional condition to the diffusion model. To this end, the generalist policy training is formulated as the standard problem of conditional generative modeling as
$\max_{\theta}\mathbb{E}_{\tau\sim\mathcal{D}}\left[\log p_{\theta}\left(x_{0}(\tau)\mid\hat{R}(\tau),\psi(l)\right)\right],$ (10)
where $\theta$ denotes diffusion parameters and $\hat{R}$ is the return-to-go. The goal is learning to generate portions of a trajectory $x_{0}(\tau)$ from the task-specific condition $\psi(l)$ via estimating the conditional data distribution with $p_{\theta}$. More details can be referred to as in Algorithms 3 and 5 in Appendix A.
@@ 4 Experiments
  • Can T2DA achieve consistent performance gain on zero-shot generalization capacity to unseen tasks? We compare it to various types of strong baselines including offline meta-RL, in-context RL, and language-conditioned policy learning approaches. (Sec. 4.1)
@@
Environments. We evaluate T2DA on three benchmarks that are widely adopted to assess generalization capacities of RL algorithms: i) the 2D navigation Point-Robot; ii) the multi-task MuJoCo locomotion control, containing Cheetah-Vel and Ant-Dir; and iii) the Meta-World manipulation platform, where a robotic arm is designed to perform a wide range of manipulation tasks, such as close faucet, lock door, open door, and press button. For each domain, we randomly sample a distribution of tasks captioned with text descriptions, and split them into training and test sets. We employ SAC [47] to independently train a single-task policy on each training task for offline dataset collection. We develop three types of offline datasets: Mixed, Medium, and Expert, following the common practice in offline RL [48]. Appendix B presents more details of environments and datasets.
Baselines. We compare T2DA to four competitive baselines that cover three representative paradigms in tackling RL generalization: the context-based offline meta-RL approaches, including 1) CSRO [9] and 2) UNICORN [26]; the in-context RL method of 3) AD [29]; and the language-conditioned policy learning method of 4) BC-Z [19]. Additionally, we include the base algorithms of T2DA implementations as two intuitive baselines: 5) DD [43] and 6) DT [44]. Appendix C gives more details.
For evaluation, all experiments are conducted using five different random seeds. The mean of received return is plotted as the bold line, with 95% bootstrapped confidence intervals of the mean indicated by the shaded region. Appendix D gives detailed implementations of T2DA, Appendix E shows tabular results for Sec. 4.2 and Sec. 4.3, and Appendix F gives analysis on fine-tuning of the text encoder.
@@
Environment T2DA-T T2DA-D BC-Z CSRO UNICORN AD DD DT
@@ 4.1 Main Results
We compare our method against various baselines under an aligned zero-shot setting. Figure 3 and Table 1 present test return curves and numerical results of converged performance on various benchmark environments using Mixed datasets. A noteworthy point is that BC-Z generally achieves better performance than offline meta-RL and in-context RL baselines, especially in harder environments like Ant-Dir and Meta-World. This again validates our motivation of exploring a much broader source of supervision from natural language, harnessing the representation power and knowledge transferability embodied in LLMs.
@@
Figure 4: Ablation results on both T2DA-D and T2DA-T using Mixed datasets. w/o world model omits pre-training the trajectory encoder, w/o alignment omits contrastive pre-training, and w/o language omits the language supervision.
Figure 5: Results of robustness to data quality where T2DA is compared against baselines using Expert and Medium datasets. Both T2DA-D and T2DA-T achieve consistent superiority across datasets of varying quality.
@@ 4.2 Ablation Study
We compare T2DA to three ablations: 1) w/o world model, it eliminates pre-training the trajectory encoder and instead updates it jointly with the text encoder during contrastive pre-training; 2) w/o alignment, it omits contrastive language-decision pre-training and fixes the text encoder initialized from pre-trained language models; and 3) w/o language, it completely removes the text encoder.
Figure 4 presents ablation results of both the T2DA-D and T2DA-T architectures. First, the elimination of pre-training the trajectory encoder (T2DA vs. w/o world model) results in decreased test returns with greater variances. It indicates that capturing environment dynamics via the world model can enable more precise and stable knowledge alignment between text and decision. Second, omitting the language-decision pre-training (w/o world model vs. w/o alignment) further decreases the performance. It confirms the existence of a semantic gap between text and decision, whereas the gap can be effectively bridged by contrastive knowledge alignment. Finally, removing the language supervision (w/o alignment vs. w/o language) can produce catastrophic performance with poor zero-shot generalization, highlighting the necessity of leveraging a wealth of knowledge from natural language. In summary, the performance of T2DA declines when any component is omitted, verifying that all components are essential to its final capabilities.
@@ 4.3 Robustness Study
Robustness to Data Quality. We compare T2DA against baselines using Expert and Medium offline datasets. As shown in Figure 5, both T2DA-D and T2DA-T consistently outperform baselines on these datasets. Notably, while most baselines suffer significant performance degradation when trained on Medium datasets, T2DA maintains high-capacity generalization despite the lower data quality. It showcases the promising applicability of T2DA in real-world scenarios where the agent often needs to learn from sub-optimal data. In summary, T2DA achieves consistent superiority across datasets of varying quality, yielding excellent robustness.
@@
Figure 6: Results of robustness to text encoders where T2DA initializes the text encoder from CLIP, BERT, and T5, showing that T2DA can efficiently harness language knowledge embedded in different language models.
@@ 4.4 Visualization Insights
We gain deep insights into the knowledge alignment between text and decision through t-SNE visualization on the example Ant-Dir task as shown in Figure 1 in Sec. 1. Appendix G shows more visualization results.
Decision Embeddings. Raw embeddings are derived from a randomly initialized trajectory encoder. In the dynamics-aware embedding space, sample points from different tasks are more clearly distinguished, and similar tasks are grouped more closely. This suggests the successful capture of task-specific environment dynamics via the world model.
Text Embeddings. Raw embeddings are derived from pre-trained language models. In the aligned space, text embeddings follow a more distinguished distribution that matches the metric of task similarity, highlighting the successful bridging of the semantic gap. Notably, the aligned embeddings follow a cyclic spectrum that matches the periodicity of angular directions in a physical sense, showing efficient comprehension of the environment dynamics.
@@ 5 Conclusions, Limitations, and Future Work
In the paper, we tackle the generalization challenge in decision domains via leveraging a much broader source of supervision from natural language. Improvements in zero-shot generalization capacities highlight the potential impact of our scalable implementations: text-to-decision diffuser and text-to-decision transformer. These findings open up new avenues for building more flexible and accessible generalist agents that can understand and act upon natural language instructions without extensive domain-specific training.
Though, our agent is trained on lightweight datasets compared to popular large models. An essential step is to implement on vast datasets with diversified domains, unleashing the scaling law with the diffusion or transformer architectures. Another step is to deploy the text-to-decision paradigm on real robots toward embodied intelligence.
@@ Appendix C Baseline Methods
  • DD [43], Decision Diffuser. It is a conditional generative modeling framework that reformulates offline decision-making through diffusion models. It reformulates RL by: 1) Modeling policies as return-conditional diffusion models that generate state trajectories; 2) Using classifier-free guidance with low-temperature sampling to extract high-quality behaviors from offline datasets; 3) Supporting flexible composition of multiple constraints or skills at test time by combining their respective noise prediction models.
  • DT [44], Decision Transformer. It is a framework that reformulates RL as sequence modeling through the transformer architecture. It has two key components: 1) Casting the RL problem as conditional sequence modeling where the model learns to generate actions by conditioning on returns and past states and actions; 2) Using a causally masked transformer architecture that enables direct credit assignment through self-attention mechanisms.
@@
Appendix D Network Architecture and Hyperparameters of T2DA
This appendix gives details of the architecture and hyperparameters of T2DA implementations as follows.
@@ Appendix D Network Architecture and Hyperparameters of T2DA
Appendix E Tabular Results for Studies of Ablation, Robustness
The results for ablation and robustness studies are presented in Table 6 and Table 7, respectively.
Table 6: Numerical results of ablation study on both T2DA-D and T2DA-T using Mixed datasets, i.e., the final test returns after learning convergence from Figure 4. w/o world model eliminates pre-training the trajectory encoder, w/o alignment omits contrastive pre-training, and w/o language removes the language supervision. Best results in bold.
Environment T2DA-T w/o world model w/o alignment w/o language T2DA-D w/o world model w/o alignment w/o language
Point-Robot 7.2
0.1
7.2
0.1
[…]
7.7
0.3
7.7
0.3
[…]
9.2
0.2
9.2
0.2
[…]
16.9
0.1
16.9
0.1
[…]
8.4
0.3
8.4
0.3
[…]
9.5
0.4
9.5
0.4
[…]
10.1
0.2
10.1
0.2
[…]
17.2
0.4
17.2
0.4
[…]
Ant-Dir 963.4
15.3
963.4
15.3
[…]
891.0
54.4
891.0
54.4
[…]
792.3
48.3
792.3
48.3
[…]
242.6
13.7
242.6
13.7
[…]
568.0
10.8
568.0
10.8
[…]
544.1
11.1
544.1
11.1
[…]
513.8
11.0
513.8
11.0
[…]
135.1
10.3
135.1
10.3
[…]
Expert T2DA-T T2DA-D BC-Z CSRO UNICORN AD DD DT
Point-Robot 6.4
0.2
6.4
0.2
[…]
7.5¯
0.2
¯7.5
0.2
[…]
14.4
0.4
14.4
0.4
[…]
12.8
0.2
12.8
0.2
[…]
11.5
1.7
11.5
1.7
[…]
17.1
0.2
17.1
0.2
[…]
17.6
0.6
17.6
0.6
[…]
17.5
0.1
17.5
0.1
[…]
Meta-World 1142.8¯
106.6
¯1142.8
106.6
[…]
1387.7
21.6
1387.7
21.6
[…]
935.5
26.7
935.5
26.7
[…]
1087.0
78.8
1087.0
78.8
[…]
787.1
16.0
787.1
16.0
[…]
991.0
37.7
991.0
37.7
[…]
913.4
5.7
913.4
5.7
[…]
1026.0
60.8
1026.0
60.8
[…]
Medium T2DA-T T2DA-D BC-Z CSRO UNICORN AD DD DT
Point-Robot 10.7
0.4
10.7
0.4
[…]
11.0¯
0.3
¯11.0
0.3
[…]
16.6
0.2
16.6
0.2
[…]
16.1
0.8
16.1
0.8
[…]
14.9
0.5
14.9
0.5
[…]
16.9
0.2
16.9
0.2
[…]
17.2
0.1
17.2
0.1
[…]
17.1
0.1
17.1
0.1
[…]
Meta-World 1033.3
130.6
1033.3
130.6
[…]
1025.3¯
43.6
¯1025.3
43.6
[…]
756.9
103.4
756.9
103.4
[…]
854.1
17.9
854.1
17.9
[…]
715.8
93.5
715.8
93.5
[…]
585.7
126.1
585.7
126.1
[…]
768.7
95.1
768.7
95.1
[…]
780.4
40.8
780.4
40.8
[…]
Appendix F Analysis of Parameter-Efficient Fine-tuning

### v3

@@ 3.4 Generalist Policy Learning
A widely adopted manner to tackle generalization in RL domains is to condition the policy on some representation that can capture task-relevant information, such as context-based meta-learning [27, 28] and prompt-based methods [10, 12]. Under this unified framework, we formulate the generalist agent as a task-conditioned policy model. There exists a true variable that represents the task identity, and we use a latent representation $h$ to approximate that variable. Then, a base policy $\pi(\cdot,h)$ can be shared across tasks by conditioning on the task representation as $\pi_{k}(a|s)\!=\!\pi(a|s;h_{k}),\forall k$. Previous studies usually approximate the task representation by acquiring expert decision data [10] or exploring the unseen task first [11]. This restricted form of supervision limits their universality and practicability since expensive high-quality samples or explorations are needed to specify any new task.
Motivated by the recent success of language models, we study a promising alternative that leverages a much broader source of supervision. Natural language is a flexible representation for transferring a variety of human ideas and intentions. Learning from its supervision can not only enhance the representation power but also enable efficient knowledge transfer. We map the textual task description $l$ to a latent embedding $\psi(l)$ via the fine-tuned text encoder in Sec. 3.3, and use it to approximate the task representation as $h\!\approx\!\psi(l)$. Then, the task-agnostic generalist policy is approximated by
$\pi(a\mid s;~{}h_{k})\approx\pi\left(a\mid s;~{}\psi(l_{k})\right),~{}~{}~{}\forall k.$ (6)
@@ 3.5 Scalale Implementations
where $c$ denotes the timestep in the forward diffusion process, and $c\!=\!0$ corresponds to the unperturbed data.  [^1] Decision diffuser diffuses over only states and trains another inverse dynamics model to generate actions. In contrast, we only keep a unified diffusion model to i) ensure simplicity with a single end-to-end model in a multi-task setting, and ii) unlock the potential of diffusion models at scale, rather than depending on additional MLPs to generate part of the information, e.g., actions. To enable a generalist policy with natural language supervision, we include the fine-tuned text embedding $\psi(l)$ as an additional condition to the diffusion model. To this end, the generalist policy training is formulated as the standard problem of conditional generative modeling as
$\max_{\theta}\mathbb{E}_{\tau\sim\mathcal{D}}\left[\log p_{\theta}\left(x_{0}(\tau)\mid\hat{R}(\tau);~{}\psi(l)\right)\right],$ (8)
where $\theta$ denotes diffusion parameters and $\hat{R}$ is the return-to-go. During evaluation, we can cast planning in RL as sampling from the diffuser. At each timestep, we observe a state $s_{t}$ in the environment, sample an $H$-step trajectory $x_{0}(\tau)$ with the diffusion process conditioned on a target return-to-go and task prompt $\psi(l)$, execute the first action $a_{t}$ in $x_{0}(\tau)$, and transition to the next state $s_{t+1}$. More details can be referred to as in Algorithms 3 and 5 in Appendix A.
Text-to-Decision Transformer (T2DA-T). We also implement T2DA with the causal transformer architecture, drawing upon the simplicity, scalability, and associated advances in language modeling such as BERT and GPT. We prepend a task prompt to the input of the causal transformer to realize generalization across tasks, akin to Prompt-DT [10]. For each task, we take the prompt-augmented trajectory $\tau^{+}$ as the input, which contains both the fine-tuned text embeddings $\psi(l)$ and the most recent $H$-step history sampled from offline datasets as
$\tau^{+}=\left(\psi(l);~{}\hat{R}_{t-H+1},s_{t-H+1},a_{t-H+1},...,\hat{R}_{t},s_{t},a_{t}\right),$ (9)
where the return-to-go $\hat{R}_{t}$ is the cumulative rewards from the current time step till the end of the episode. Then, the model predicts actions autoregressively using a causal self-attention mask. This architecture introduces only one additional token compared to the standard decision transformer, allowing minimal architecture change and lightweight cost for generalist training. More details can be referred to as in Algorithms 4 and 6 in Appendix A.
@@ 4 Experiments
  • Can T2DA achieve consistent performance gain on zero-shot generalization capacity to unseen tasks? We compare it to various types of strong baselines, including offline meta-RL, in-context RL, and language-conditioned policy learning approaches. (Sec. 4.1)
@@
Environments. We evaluate T2DA on three benchmarks that are widely adopted to assess generalization capacities of RL algorithms: i) the 2D navigation Point-Robot; ii) the multi-task MuJoCo locomotion control, containing Cheetah-Vel and Ant-Dir; and iii) the Meta-World platform for robotic manipulation. For each domain, we randomly sample a distribution of tasks captioned with text descriptions, and split them into training and test sets. We employ SAC [52] to independently train a single-task policy on each training task for offline dataset collection. We develop three types of offline datasets: Mixed, Medium, and Expert, following the common practice in offline RL [53]. Appendix B presents more details of environments and datasets.
Baselines. We compare T2DA to five competitive baselines that cover three representative paradigms in tackling RL generalization: the context-based offline meta-RL approaches, including 1) CSRO [11] and 2) UNICORN [29]; the in-context RL method of 3) AD [32]; and the language-conditioned policy learning methods including 4) BC-Z [21] and 5) BAKU [44].
For evaluation, all experiments are conducted using five different random seeds. The mean of received return is plotted as the bold line, with 95% bootstrapped confidence intervals of the mean indicated by the shaded region. Appendix D gives detailed implementations of T2DA, Appendix LABEL:app:tab shows tabular results for Sec. 4.3, and Appendix E gives analysis on fine-tuning of the text encoder.
@@
Environment T2DA-T T2DA-D BC-Z CSRO UNICORN AD BAKU
@@ 4.1 Main Results
We compare our method against various baselines under an aligned zero-shot setting. Figure 3 and Table 1 present test return curves and numerical results of converged performance on various benchmark environments using Mixed datasets. A noteworthy point is that language-conditioned baselines such as BC-Z and BAKU generally achieve better performance than offline meta-RL and in-context RL baselines, especially in harder environments like Ant-Dir and Meta-World. This again validates our motivation of exploring a much broader source of supervision from natural language, harnessing the representation power and knowledge transferability embodied in pre-trained LLMs.
@@
Figure 4: Ablation results using Mixed datasets. w/o world omits pre-training the trajectory encoder, w/o align omits contrastive pre-training, and w/o text omits the language supervision.
Table 2: Numerical results of ablation study on both T2DA-D and T2DA-T using Mixed datasets, i.e., the final test returns after learning convergence from Figure 4. Best results in bold.
Environment T2DA-T w/o world w/o align w/o text T2DA-D w/o world w/o align w/o text
Point-Robot 7.2
0.1
7.2
0.1
[…]
7.7
0.3
7.7
0.3
[…]
9.2
0.2
9.2
0.2
[…]
16.9
0.1
16.9
0.1
[…]
8.4
0.3
8.4
0.3
[…]
9.5
0.4
9.5
0.4
[…]
10.1
0.2
10.1
0.2
[…]
17.2
0.4
17.2
0.4
[…]
Ant-Dir 963.4
15.3
963.4
15.3
[…]
891.0
54.4
891.0
54.4
[…]
792.3
48.3
792.3
48.3
[…]
242.6
13.7
242.6
13.7
[…]
568.0
10.8
568.0
10.8
[…]
544.1
11.1
544.1
11.1
[…]
513.8
11.0
513.8
11.0
[…]
135.1
10.3
135.1
10.3
[…]
Figure 5: Results of robustness to data quality, where T2DA is compared to baselines using Expert and Medium datasets. T2DA-D and T2DA-T achieve consistent superiority across various datasets.
Expert T2DA-T T2DA-D BC-Z CSRO UNICORN AD BAKU
Point-Robot 6.4
0.2
6.4
0.2
[…]
7.5¯
0.2
¯7.5
0.2
[…]
14.4
0.4
14.4
0.4
[…]
12.8
0.2
12.8
0.2
[…]
11.5
1.7
11.5
1.7
[…]
17.1
0.2
17.1
0.2
[…]
15.4
0.1
15.4
0.1
[…]
Meta-World 1142.8¯
106.6
¯1142.8
106.6
[…]
1387.7
21.6
1387.7
21.6
[…]
935.5
26.7
935.5
26.7
[…]
1087.0
78.8
1087.0
78.8
[…]
787.1
16.0
787.1
16.0
[…]
991.0
37.7
991.0
37.7
[…]
833.6
8.3
833.6
8.3
[…]
Medium T2DA-T T2DA-D BC-Z CSRO UNICORN AD BAKU
Point-Robot 10.7
0.4
10.7
0.4
[…]
11.0¯
0.3
¯11.0
0.3
[…]
16.6
0.2
16.6
0.2
[…]
16.1
0.8
16.1
0.8
[…]
14.9
0.5
14.9
0.5
[…]
16.9
0.2
16.9
0.2
[…]
16.1
0.1
16.1
0.1
[…]
Meta-World 1033.3
130.6
1033.3
130.6
[…]
1025.3¯
43.6
¯1025.3
43.6
[…]
756.9
103.4
756.9
103.4
[…]
854.1
17.9
854.1
17.9
[…]
715.8
93.5
715.8
93.5
[…]
585.7
126.1
585.7
126.1
[…]
669.5
11.4
669.5
11.4
[…]
@@ 4.2 Ablation Study
We compare T2DA to three ablations: 1) w/o world, it eliminates pre-training the trajectory encoder and instead updates it jointly with the text encoder during contrastive pre-training; 2) w/o align, it omits contrastive language-decision pre-training and fixes the text encoder initialized from pre-trained language models; and 3) w/o text, it completely removes the text encoder.
Figure 4 presents ablation results of both architectures. First, the erasure of pre-training the trajectory encoder (T2DA vs. w/o world) results in decreased test returns with greater variances. It indicates that capturing environment dynamics via the world model can enable more precise and stable knowledge alignment between text and decision. Second, omitting the language-decision pre-training (w/o world vs. w/o align) further decreases the performance. It confirms the existence of a semantic gap between text and decision, whereas the gap can be effectively bridged by contrastive knowledge alignment. Finally, removing the language supervision (w/o align vs. w/o text) can produce catastrophic performance with poor zero-shot generalization, highlighting the necessity of leveraging a wealth of knowledge from natural language. In summary, the performance of T2DA declines when any component is omitted, verifying that all components are essential to its final capabilities.
@@ 4.3 Robustness Study
Robustness to Data Quality. We compare T2DA against baselines using Expert and Medium offline datasets. As shown in Figure 5, both T2DA-D and T2DA-T consistently outperform baselines on these datasets. Notably, while most baselines suffer significant performance degradation when trained on Medium datasets, T2DA maintains high-capacity generalization despite the lower data quality.
Figure 6: Results of robustness to text encoders where T2DA initializes the text encoder from CLIP, BERT, and T5.
It showcases the promising applicability of T2DA in real-world scenarios where the agent often needs to learn from sub-optimal data. In summary, T2DA achieves consistent superiority across datasets of varying quality, yielding excellent robustness.
@@ 4.4 Visualization Insights
Figure 7: t-SNE visualization of Cheetah-Vel where tasks with target velocities in $[0.075,3.0]$ are mapped into rainbow-colored points. This inspiring finding verifies the effective alignment of text embeddings to comprehend environment dynamics and interpretable language grounding.
We gain deep insights into the knowledge alignment between text and decision through t-SNE visualization on the example Cheetah-Vel task as shown in Figure 7.
Decision Embeddings. Raw decision embeddings are derived from a randomly initialized trajectory encoder. In the dynamics-aware embedding space, sample points from different tasks are more clearly distinguished, and similar tasks are grouped more closely. The initially entangled trajectories are transformed into well-separated clusters, highlighting the successful capture of task-specific environment dynamics via the generalized world model.
Text Embeddings. Raw text embeddings are derived from pre-trained language models, and are scattered with a vague distribution structure. In the aligned space, text embeddings follow a more distinguished distribution that matches the metric of task similarity, highlighting successful bridging of the semantic gap. Notably, the aligned embeddings form a clear rectilinear distribution from red (low velocity) to blue (high velocity), exactly matching the rectilinear spectrum of target velocities in a physical sense. It verifies efficient comprehension of the environment dynamics.
@@ 5 Conclusions, Limitations, and Future Work
In the paper, we tackle offline meta-RL challenge via leveraging a much broader source of supervision from natural language. Improvements in zero-shot generalization capacities highlight the potential impact of our scalable implementations: text-to-decision diffuser and text-to-decision transformer. These findings open up new avenues for building more flexible and accessible generalist agents that can understand and act upon natural language instructions without extensive domain-specific training.
Though, our agent is trained on lightweight datasets compared to popular large models. An essential step is to implement on vast datasets with diversified domains, unleashing the scaling law with the diffusion or transformer architectures. Another step is to deploy the idea of knowledge alignment in VLA models, or deploy the text-to-decision paradigm on real robots toward embodied intelligence.
@@ B.1 The Details of Environments
      handle-pull-side       Pull a handle up sideways
      hand-insert       Insert the gripper into a hole
@@ Appendix C Baseline Methods
  • BAKU [44], an efficient transformer architecture for multi-task policy learning that improves upon prior work in offline imitation learning. It carefully integrates multiple key components: observation trunks for multi-modal fusion, action chunking for smoother control, multi-sensory observation processing, and modular action heads for flexible prediction. The model is optimized through behavior cloning with the objective: $\min\sum_{\text{task}~{}k}\sum_{(o,a)\sim\mathcal{D}}|a-\pi(a|o,T_{k})|^{2}$, where $o$ represents multi-modal observations and $T_{k}$ denotes the task instruction. For a fair comparison, we adjust BAKU to the language-conditioned zero-shot setting.
@@
Appendix D Implementation Details of T2DA
D.1 Network Architecture and Hyperparameters
@@
D.2 Computation
We train our models on one Nvidia RTX4080 GPU with the Intel Core i9-10900X CPU and 256G RAM. Dynamic-aware decision embedding and constrastive language-decision pre-training cost about 0.5 hours. The generalist policy training cost about 0.5-3 hours depending on the complexity of the environment.
Appendix E Analysis of Parameter-Efficient Fine-tuning